# Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime

### v1

@@ 2.4 Provenance and transport of magnetic fields
The presence and morphology of magnetic field structures are inextricably linked to the evolution and accretion dynamics of protoplanetary discs. Yet, due to current observational limitations, a complete picture for the magnetic field morphology remains elusive (Tsukamoto et al., 2023). Existing measurements tentatively suggest a weak vertical field dominated by a toroidal component, consistent with the configuration expected under the weak magnetic pressure wind accretion paradigm in the dead zone. For instance, the circular polarisation of Zeeman-split CN molecular lines place tentative upper limits on the vertical magnetic field strength of $B_{z}<0.8$ mG in TW Hya (Vlemmings et al., 2019) and $B_{z}<6.1$ mG in AS 209 (Harrison et al., 2021). Meanwhile, polarised thermal emission from aligned non-spherical dust grains implies a total magnetic field strength of $B\sim 0.3$ mG in HD 142527, with the toroidal component dominating by a factor of four (Ohashi et al., 2025), although this interpretation may be affected by non-magnetic grain alignment mechanisms (see Harrison et al., 2024).
@@
The leading theory for the origin of a poloidal magnetic field threading the disc is inheritance from the protostellar nebulae (e.g. Kunz & Mouschovias, 2010; Crutcher, 2012), where magnetic flux is drawn inward during core collapse, shaping the field into a pinched, hourglass configuration. However, recent simulations of MRI turbulent accretion discs by Jacquemin-Ide et al. (2024), suggest that large-scale poloidal fields can also be generated in situ via the MRI dynamo.
@@
Nevertheless, it is undoubtedly the case that the transport of such a large-scale magnetic field impacts disc evolution. The conventional advection-diffusion paradigm sets this magnetic flux transport as a competition between inward advection, due to bulk accretion flow, and outward diffusion, driven by turbulent or physical resistivity (e.g. Lubow et al., 1994; Heyvaerts et al., 1996). Whilst extensive (semi-)analytic work has highlighted the importance of vertical disc structure (Rothstein & Lovelace, 2008; Guilet & Ogilvie, 2012, 2013) and non-ideal MHD effects (e.g. Leung & Ogilvie, 2019), a complete local model of flux transport remains elusive. As such, global numerical simulations are essential.
@@
In protoplanetary discs, global simulations have revealed an intriguing paradigm for magnetic flux transport. In the ideal MHD regime, recent studies suggest that advection dominates and flux is transported inward (Zhu & Stone, 2018; Jacquemin-Ide et al., 2021). Meanwhile, in the laminar, non-ideal MHD regime, diffusion tends to dominate and typically drives outward flux transport (e.g Bai, 2017; Gressel et al., 2020; Lesur, 2021b), though this picture is complicated by the Hall effect (Bai & Stone, 2017). Therefore, flux transport appears to diverge around the dead–active zone interface, and the fate of the inner magnetic field as it interacts with the stellar field remains an open question. As such, the global evolution of magnetic flux is expected to be strongly dependent on the interactions around the dead–active zone interface, with the star–disc boundary (e.g. Pantolmos et al., 2020; Takasao et al., 2025; Zhu, 2025) and the initial magnetic field configuration.
@@ 3 Model and Numerical Methods
The global three-dimensional non-ideal MHD ZNF simulations in this study depend on the governing equations outlined in Section 3.1 and the inner disc model detailed in Section 3.2. The corresponding numerical methods are presented in Section 3.3, followed by the key diagnostics in Section 3.4.
@@ 3.1 Governing equations
The disc model is evolved using the non-ideal MHD equations, in which the Hall effect is neglected. These determine the evolution of the density $\rho$, velocity field $\mathbf{u}$, magnetic field $\mathbf{B}$ and total energy density $E=e+\rho u^{2}/2+B^{2}/8\pi$, where $e$ is the internal energy density, through,
@@
$\partial_{t}\mathbf{B}$ =×$=-\nabla\times\left(\mathbf{\mathcal{E}_{I}}+\mathbf{\mathcal{E}_{NI}}\right)$ (3)
@@
where the ideal and non-ideal components of the electromotive field are respectively given by,
$\mathbf{\mathcal{E}_{I}}=\mathbf{u}\times\mathbf{B}\;\;\textrm{and}\;\;\mathbf{\mathcal{E}_{NI}}=\frac{4\pi}{c}\left[-\eta_{O}\mathbf{J}+\eta_{A}(\mathbf{J}\times\mathbf{b})\times\mathbf{b}\right],$ (5)
under the Gaussian unit convention. Here, $\Phi=-GM/r$ is the gravitational potential of a central object of mass $M$ in spherical coordinates $(r,\theta,\phi)$, $P$ the gas pressure, $c$ the speed of light, $\mathbf{J}=(c/4\pi)\mathbf{\nabla}\times\mathbf{B}$ the current density, $\mathcal{L}$ an imposed energy density thermal relaxation term, $\mathbf{b}$ the direction of the magnetic field, $\eta_{O}$ the Ohmic diffusivity and $\eta_{A}$ the ambipolar diffusivity.
The system of equations is closed using the equation of state for a perfect gas, $P=e(\gamma-1)$, where $\gamma$ is the adiabatic index. To model disc thermodynamics, we adopt a $\beta$-cooling prescription model and relax the system toward a target temperature profile $T_{\textrm{eff}}(R,z)$ using,
@@
We set $\gamma=1.05$ and adapt a characteristic thermal relaxation timescale of $\tau=0.1\Omega_{K}^{-1}$, such that the vertical shear instability is marginal (Nelson et al., 2013), where $\Omega_{K}^{2}=GMR^{-3}$ is the squared Keplerian angular frequency. This rapid readjustment leads to an effective quasi-locally isothermal equation of state, $P=c_{s}^{2}\rho$, where the local isothermal sound speed in the disc is $c_{s}=H\Omega_{K}$, for the disc scale height $H$.
@@
3.2 Inner disc model
The inner protoplanetary disc model introduced in this section is characterised by the following key features: (i) it is magnetised with an MRI dead–active zone interface at a static cylindrical radius which is modelled by prescribed profiles of Ohmic diffusion and ambipolar diffusion; (ii) it is cool and embedded in a hot, low-density corona; (iii) it is dynamically isolated from the star such that it only feels the central gravitational potential; (iv) it neglects both self-gravity and explicit viscosity; and (v) it is dust free.
@@ 3.2.1 Temperature prescription
$T_{\textrm{eff}}(R,z)=\frac{1}{2}\left[\left(T_{\textrm{disc}}+T_{\textrm{cor}}\right)+\left(T_{\textrm{disc}}-T_{\textrm{cor}}\right)\tanh\left(\frac{\left|z\right|-4\varepsilon R_{i}}{R_{0}}\right)\right],$ (7)
is exhibited in Fig. 1. The model splits the domain into a cold dense disc with temperature $T_{\textrm{disc}}(R)=T_{0}R_{0}\varepsilon^{2}/R_{i}$ and a hot, low-density corona with temperature $T_{\textrm{cor}}(R)=16T_{\textrm{disc}}(R_{i})$, where $\varepsilon=H/R$ is the disc aspect ratio. The subscript 00 refers to the midplane at the inner edge of the domain ($R=R_{\textrm{in}},z=0$), and $R_{i}=\max(R,R_{0})$.
The simplified temperature profile prescribes a hot corona to account for the mixture of stellar irradiation (e.g. Aresu et al., 2011) and the inefficient gas cooling resulting from reduced thermal accommodation on dust particles in this dust-sparse region (see fig. 10 in Thi et al., 2019). The transition height is fixed at $z=\pm 4H$, following the ionisation model of the inner disc presented in fig. 1 of Lesur et al. (2014).
@@ 3.2.2 Non-ideal MHD model and the dead–active zone interface
To capture the complex dynamic and thermodynamic behaviour of the dead–active zone interface, a broad range of numerical models have been used in previous work. These include: (i) spatially static HD viscosity transitions (e.g. Lyra et al., 2009; Miranda et al., 2016; Ueda et al., 2019); (ii) spatially static Ohmic resistivity transitions (e.g. Dzyurkevich et al., 2010; Okuzumi & Hirose, 2011; Lyra & Mac Low, 2012), sometimes extended to include ambipolar diffusion (Iwasaki et al., 2024); and (iii) temperature-dependent transitions in either viscosity (e.g. Ueda et al., 2019; Cecil & Flock, 2024) or Ohmic resistivity (e.g. Faure et al., 2014, 2015; Faure & Nelson, 2016; Flock et al., 2017).
@@
$\text{R}_{m}=\frac{\Omega_{K}H^{2}}{\eta_{O}}\quad\text{and}\quad\Lambda_{A}=\frac{v^{2}_{\text{A}}}{\Omega_{K}\eta_{A}},$ (8)
@@
$\Lambda_{A}(R,z)$ $=\frac{\Lambda_{A_{0}}}{4}\left[1-\tanh\left(\frac{|z|-4\varepsilon R}{0.2R_{0}}\right)\right]\left[1-\tanh\left(\frac{R-R_{\textrm{DZI}}}{0.02R_{0}}\right)\right],$ (9)
$\text{R}_{m}(R,z)$ $=\frac{\text{R}_{m_{0}}\rho_{0}R_{0}}{4\rho R}\left[1-\tanh\left(\frac{|z|-4\varepsilon R}{0.2R_{0}}\right)\right]\left[1-\tanh\left(\frac{R-R_{\textrm{DZI}}}{0.02R_{0}}\right)\right],$
as shown in Fig. 1, with $\Lambda_{A_{0}}=1$ and $\text{R}_{m_{0}}=25$ guided by estimates from fig. 8 in Thi et al. (2019) and fig. 10 in Lesur (2021a). The $\text{R}_{m}$ scaling with $\rho$ and $R$ follows Lesur (2021b), assuming constant $\Lambda_{A}$, a two-species plasma (electrons and ions), and dominant ions, primarily HCO+, that are much heavier than the main neutral component, H2.
Therefore, the model mimics that ionisation in the surface layers becomes effective at the same height as the disc–corona temperature transition, $|z|\gtrsim 4H$, and imposes a static cylindrical dead–active zone interface at ${R_{\textrm{DZI}}}=4R_{0}$. This interface is sharply defined to reflect the steep radial variation in the ionisation fraction in the inner disc, whether driven by thermal ionisation of alkali metals (Na, K, etc.) via the Saha equation or by thermionic and ion emission from dust grains (Desch & Turner, 2015).
@@ 3.3 Numerical methods
The global non-ideal MHD simulations presented in this work are performed using the idefix code, which integrates the compressible MHD equations using a finite-volume high-order Godunov method. The interpolation to cell interfaces is performed using a linear piecewise reconstruction scheme with a van Leer slope limiter, and inter cell-fluxes are computed using the HLLD approximate Riemann solver (for motivation see Flock et al., 2011). The quantities are evolved in time using a total-variation diminishing second order Runge-Kutta method with the Courant number set to $C=0.5$. The solenoidal condition ($\mathbf{\nabla}\cdot\mathbf{B}=0$) is enforced to machine precision via the constrained transport method (Evans & Hawley, 1988) and an electromotive field reconstruction based on the $\mathcal{E}^{c}$ scheme (Gardiner & Stone, 2005). Parabolic non-ideal MHD diffusivity terms are handled by a centred finite-difference method. All simulations were run on NVIDIA A100 GPUs using double precision arithmetic.
In total, we present five production simulations: a fiducial simulation, ZNF-FID, and four extensions, each designed to test the robustness of the results by varying the the numerical domain or the initial conditions (see Table 3.3 for grid details). ZNF-HRES serves as a resolution test. ZNF-LONG explores longer-term evolution. ZNF-AZ removes the non-ideal MHD diffusivity prescription, such that the disc is fully MRI turbulent. ZNF-MRI-IC, begins from an initial condition of a fully MRI turbulent disc, onto which the non-ideal MHD diffusivity profile is imposed at $t_{\text{in}}=0$.
@@
tableNumerical properties of the five three-dimensional ZNF global MHD simulations. Each is integrated up to $t_{\textrm{in}}^{\textrm{end}}$ on a domain with azimuthal extent, $\phi_{\textrm{max}}$. ZNF-MRI-IC refers to a simulation where the non-ideal MHD profile is imposed on top of a fully MRI turbulent disc at $t_{\textrm{in}}=0$, while ZNF-AZ refers to a simulation where the disc is in the ideal MHD regime, such that there is no exterior dead zone, or dead–active zone interface. The total computational cost of these production simulations was approximately $3000$ GPU hours. Model name $\left(N_{r},N_{\theta},N_{\phi}\right)$ $\phi_{\textrm{max}}$ $t_{\text{in}}^{\text{end}}$ Reason ZNF-FID $(512,192,512)$ $2\pi$ 750 fiducial ZNF-HRES $(1024,192,1024)$ $2\pi$ 375 resolution ZNF-LONG $(512,192,256)$ $\pi$ 1800 late time ZNF-MRI-IC $(512,192,256)$ $\pi$ 750 initial condition ZNF-AZ $(512,192,256)$ $\pi$ 750 full active

### v3

@@ 2.4 Provenance and transport of magnetic fields
The presence and morphology of magnetic field structures are linked to the evolution and accretion dynamics of protoplanetary discs. Yet, due to current observational limitations, a complete picture for the magnetic field morphology remains elusive (tsukamoto_role_2023). Existing measurements tentatively suggest a weak vertical field dominated by a toroidal component, consistent with the configuration expected under the weak magnetic-pressure wind accretion paradigm in the dead zone. For instance, the circular polarisation of Zeeman-split CN molecular lines places tentative upper limits on the vertical magnetic field strength of $B_{z}<0.8$ mG in TW Hya (vlemmings_stringent_2019) and $B_{z}<6.1$ mG in AS 209 (harrison_alma_2021). Meanwhile, polarised thermal emission from aligned non-spherical dust grains implies a total magnetic field strength of $B\!\sim\!0.3$ mG in HD 142527, with the toroidal component dominating by a factor of four (ohashi_observationally_2025), although non-magnetic grain alignment mechanisms (see harrison_protoplanetary_2024) endanger this interpretation.
@@
The leading theory for the origin of a poloidal magnetic field threading the disc is inheritance from the protostellar nebulae (e.g. kunz_nonisothermal_2010; crutcher_magnetic_2012), where magnetic flux is drawn inward during core collapse, shaping the field into a pinched, hourglass configuration. However, recent simulations of MRI-turbulent accretion discs by jacquemin-ide_magnetorotational_2024, suggest that large-scale poloidal fields can also be generated in situ via the MRI dynamo.
@@
Nevertheless, it is undoubtedly the case that the transport of such a large-scale magnetic field impacts disc evolution. The conventional advection-diffusion paradigm sets this magnetic flux transport as a competition between inward advection, due to bulk accretion flow, and outward diffusion, driven by turbulent or physical resistivity (e.g. lubow_magnetic_1994; heyvaerts_magnetic_1996). Whilst extensive (semi-)analytic work has highlighted the importance of vertical disc structure (rothstein_advection_2008; guilet_transport_2012; guilet_transport_2013) and non-ideal MHD effects (e.g. leung_local_2019), a complete local model of flux transport remains elusive. As such, global numerical simulations are essential.
@@
Global simulations suggest an intriguing paradigm for magnetic flux transport in the inner regions of protoplanetary discs. In the ideal-MHD regime, recent studies show that advection dominates and flux is transported inward (zhu_global_2018; jacquemin-ide_magnetic_2021). Meanwhile, in the laminar, non-ideal-MHD regime, diffusion tends to dominate and typically drives flux outward (e.g bai_global_2017; gressel_global_2020; lesur_systematic_2021), though this picture is complicated by the Hall effect (bai_hall_2017). Therefore, the transport of flux transport is divergent around the dead–active zone interface. Moreover, the fate of the inner magnetic field as it drifts inward and interacts with the stellar field remains an open question. As such, the global evolution of magnetic flux is expected to be strongly dependent on the interactions around the dead–active zone interface, with the star–disc boundary (e.g. pantolmos_magnetic_2020; takasao_connecting_2025; zhu_global_2025), and the initial magnetic field configuration.
@@ 3 Model and Numerical Methods
The global three-dimensional non-ideal MHD ZNF simulations in this study evolve the governing equations outlined in Section 3.1 as applied to the inner disc model detailed in Section 3.2. The corresponding numerical methods are presented in Section 3.3, followed by the key diagnostics in Section 3.4.
@@ 3.1 Governing equations
The evolution of our disc model is governed by the non-ideal MHD equations, in which Ohmic and ambipolar diffusion are included but the Hall effect neglected. These determine the evolution of the density $\rho$, velocity field $\mathbf{u}$, magnetic field $\mathbf{B}$, and energy density $E=e+\rho u^{2}/2+B^{2}/8\pi$, where $e$ is the internal energy density, through,
@@
$\partial_{t}\mathbf{B}$ $=-\nabla\times\left(\mathbf{\mathcal{E}_{I}}+\mathbf{\mathcal{E}_{NI}}\right)$ (3)
@@
where the ideal and non-ideal components of the electromotive field are given by, respectively,
$\mathbf{\mathcal{E}_{I}}=\mathbf{u}\times\mathbf{B}\;\;\textrm{and}\;\;\mathbf{\mathcal{E}_{NI}}=\frac{4\pi}{c}\left[-\eta_{\text{O}}\mathbf{J}+\eta_{\text{A}}(\mathbf{J}\times\mathbf{b})\times\mathbf{b}\right],$ (5)
under the Gaussian unit convention. Here, $\Phi=-GM/r$ is the gravitational potential of a central spherical star of mass $M$ in spherical coordinates $(r,\theta,\phi)$, $P$ the gas pressure, $c$ the speed of light, $\mathbf{J}=(c/4\pi)\mathbf{\nabla}\times\mathbf{B}$ the current density, $\mathcal{L}$ an imposed energy-density thermal-relaxation term, $\mathbf{b}$ the direction of the magnetic field, $\eta_{O}$ the Ohmic diffusivity, and $\eta_{\text{A}}$ the ambipolar diffusivity.
The system of equations is closed using the equation of state for a perfect gas, $P=e(\gamma-1)$, where $\gamma$ is the adiabatic index. To ensure that the integration scheme is positive-definite with respect to $P$, particularly in the low-density regions, we adopt $\gamma=1.05$. The disc temperature is controlled through a $\beta$-cooling prescription: the system is relaxed towards a target temperature, $T_{\textrm{eff}}(R,z)$, using,
@@
The characteristic thermal-relaxation timescale is $\tau=0.1\Omega_{\text{K}}^{-1}$, so that the vertical shear instability is marginal (nelson_linear_2013), where $\Omega_{\text{K}}^{2}=GMR^{-3}$ is the squared Keplerian angular frequency. This rapid thermal readjustment leads to an effectively quasi-local isothermal equation of state, $P=c_{\text{s}}^{2}\rho$, where the local isothermal sound speed in the disc is $c_{\text{s}}=H\Omega_{\text{K}}$, for the disc scale height $H$.
@@
3.2 Inner-disc model
Our model of the inner disc regions is characterised by the following features: (i) it is magnetised with a dead–active zone interface at a fixed cylindrical radius, modelled using prescribed profiles of Ohmic diffusion and ambipolar diffusion; (ii) it is cool and embedded in a hot, low-density corona; (iii) apart from the central gravitational potential, the disc is dynamically isolated from the star; (iv) it neglects both self-gravity and explicit viscosity; and (v) it is dust-free.
@@ 3.2.1 Temperature prescription
$T_{\textrm{eff}}(R,z)=\frac{1}{2}\left[\left(T_{\textrm{disc}}+T_{\textrm{cor}}\right)+\left(T_{\textrm{disc}}-T_{\textrm{cor}}\right)\tanh\left(\frac{\left|z\right|-4\varepsilon R_{\text{i}}}{R_{0}}\right)\right],$ (7)
is exhibited in Fig. 1. The model splits the domain into a cold dense disc with temperature $T_{\textrm{disc}}(R)=T_{0}R_{0}\varepsilon^{2}/R_{\text{i}}$, and a hot, low-density corona with temperature $T_{\textrm{cor}}(R)=16T_{\textrm{disc}}(R)$. Here $\varepsilon=H/R$ is the disc aspect ratio, the subscript 0 refers to the midplane at the inner edge of the domain, and $R_{\text{i}}=\max(R,R_{0})$, which ensures consistency with future models that are closed at the poles. Thus the radial profile of the effective temperature scales as $T\propto R^{-1}$ when $R\geq R_{0}$. This enforces a constant aspect ratio. We adopt $\varepsilon=0.1$ as a compromise between numerical efficiency – ensuring that the MRI is adequately resolved – and the expected inner disc thickness of roughly 0.03–0.05 at $R=1$ au (e.g. bitsch_structure_2015).
This simplified temperature profile prescribes a hot corona to account for the mixture of stellar irradiation (e.g. aresu_x-ray_2011) and the inefficient gas cooling resulting from reduced thermal accommodation on dust particles in this dust-sparse region (see fig. 10 in thi_radiation_2019). The transition height is fixed at $z=\pm 4H$, following the ionisation model presented in fig. 1 of lesur_thanatology_2014.
@@ 3.2.2 Non-ideal MHD model and the dead–active zone interface
To capture the complex dynamic and thermodynamic behaviour of the dead–active zone interface, a broad range of numerical models have been used in previous work. These include: (i) spatially static HD viscosity transitions (e.g. lyra_planet_2009; miranda_rossby_2016; ueda_dust_2019); (ii) spatially static Ohmic resistivity transitions (e.g. dzyurkevich_trapping_2010; okuzumi_modeling_2011; lyra_rossby_2012), sometimes extended to include ambipolar diffusion (iwasaki_dynamics_2024); and (iii) temperature-dependent transitions in either viscosity (e.g. ueda_dust_2019; cecil_variability_2024) or Ohmic resistivity (e.g. faure_thermodynamics_2014; faure_vortex_2015; faure_planet_2016; flock_3d_2017).
@@
$\text{R}_{\text{m}}=\frac{\Omega_{\text{K}}H^{2}}{\eta_{\text{O}}}\quad\text{and}\quad\Lambda_{\text{A}}=\frac{v^{2}_{\text{A}}}{\Omega_{\text{K}}\eta_{\text{A}}},$ (8)
@@
$\text{R}_{\text{m}}(R,z)=$ $\,\frac{4\text{R}_{\text{m}_{0}}\rho_{0}R_{0}}{\rho R}\left[1-\tanh\left(\frac{|z|-4\varepsilon R}{0.2R_{0}}\right)\right]^{-1}$ (9)
$\times\left[1+\tanh\left(\frac{R-R_{\textrm{DZI}}}{0.02R_{0}}\right)\right]^{-1},$
$\Lambda_{\text{A}}(R,z)=$  $\,4\Lambda_{\text{A}_{0}}\left[1-\tanh\left(\frac{|z|-4\varepsilon R}{0.2R_{0}}\right)\right]^{-1}\left[1+\tanh\left(\frac{R-R_{\textrm{DZI}}}{0.02R_{0}}\right)\right]^{-1},$
as shown in Fig. 1, with $\Lambda_{\text{A}_{0}}=1$ and $\text{R}_{\text{m}_{0}}=25$ guided by estimates from fig. 8 in thi_radiation_2019 and fig. 10 in lesur_magnetohydrodynamics_2021. We take the dead–active zone interface to be at ${R_{\textrm{DZI}}}=4R_{0}$. The $\text{R}_{\text{m}}$ scaling with $\rho$ and $R$ follows lesur_systematic_2021, assuming constant $\Lambda_{\text{A}}$, a two-species plasma (electrons and ions), and dominant ions, primarily HCO+, that are much heavier than the main neutral component, H2.
Therefore, the model assumes that ionisation in the surface layers becomes effective at the same height as the disc–corona temperature transition, $|z|\gtrsim 4H$, and it imposes a static cylindrical dead-zone interface, $R=R_{\text{DZI}}$. This interface is sharply defined to reflect the steep radial variation in the ionisation fraction in the inner disc, whether driven by thermal ionisation of alkali metals (Na, K, etc.) via the Saha equation or by thermionic and ion emission from dust grains (desch_high-temperature_2015; williams_ionization_2024).
@@ 3.3 Numerical methods
The global non-ideal MHD simulations presented in this work are performed using the idefix code, which integrates the compressible MHD equations using a finite-volume high-order Godunov method. The interpolation to cell interfaces is performed using a linear piecewise reconstruction scheme with a van Leer slope limiter, and inter cell-fluxes are computed using the HLLD approximate Riemann solver (for motivation see flock_turbulence_2011). The quantities are evolved in time using a total-variation diminishing second order Runge–Kutta method with the Courant number set to $\text{C}=0.5$. The solenoidal condition ($\mathbf{\nabla}\cdot\mathbf{B}=0$) is enforced to machine precision via the constrained transport method (evans_simulation_1988) and an electromotive-field reconstruction based on the $\mathcal{E}^{\text{c}}$ scheme (gardiner_unsplit_2005). Parabolic non-ideal MHD diffusivity terms are handled by a centred finite-difference method. All simulations were run on NVIDIA A100 GPUs using double-precision arithmetic. [^2] ZNF-FID used roughly 800 GPU-hours to perform $6.2\times 10^{6}$ integration cycles, achieving $1.8\times 10^{9}$ cell updates in total per second across 16 GPUs, using a 4–1–4 decomposition in $(r,\theta,\phi)$.
@@
tableNumerical properties of the five three-dimensional ZNF global MHD simulations. Each is integrated up to $t_{\textrm{in}}^{\textrm{end}}$ on a domain with azimuthal extent $\phi_{\textrm{max}}$. ZNF-MRI-IC refers to a simulation where the non-ideal MHD profile is imposed on top of a fully MRI-turbulent disc at $t_{\textrm{in}}=0$, whilst ZNF-AZ refers to a simulation where the disc is in the ideal-MHD regime, such that there is no exterior dead zone, or dead–active zone interface. The total computational cost of these five production simulations was approximately $3000$ GPU hours. Model name $\left(N_{r},N_{\theta},N_{\phi}\right)$ $\phi_{\textrm{max}}$ $t_{\text{in}}^{\text{end}}$ Reason ZNF-FID $(512,192,512)$ $2\pi$ 750 fiducial ZNF-HRES $(1024,192,1024)$ $2\pi$ 375 resolution ZNF-LONG $(512,192,256)$ $\pi$ 1800 late time ZNF-MRI-IC $(512,192,256)$ $\pi$ 750 initial condition ZNF-AZ $(512,192,256)$ $\pi$ 750 full active
In total, we present five production simulations: a fiducial simulation, ZNF-FID, and four extensions, each designed to test the robustness of the results by varying the numerical domain or the initial conditions (see Table 3.3 for grid details). ZNF-HRES serves as a resolution test. ZNF-LONG explores longer-term evolution. ZNF-AZ removes the non-ideal MHD diffusivity prescription, such that the disc is fully MRI turbulent. ZNF-MRI-IC, begins from a fully MRI-turbulent initial condition, onto which the non-ideal MHD diffusivity profile is imposed at $t_{\text{in}}=0$.